OpenStudy (anonymous):

There are 4 boxes colored red, yellow, green and blue. If 2 boxes are selected, how many combinations are there for at least one green box or one red box to be selected? a..1 b.. 6 c.. 9 d… 5

OpenStudy (anonymous):

come on gys

OpenStudy (anonymous):

c

OpenStudy (anonymous):

wait wait

OpenStudy (anonymous):

Its 5

OpenStudy (anonymous):

my answer is 5

OpenStudy (anonymous):

how u solve it

OpenStudy (anonymous):

select green first..then there are 3 ways to select the second box..you get 3 ways if you select red box first..but there is one combination repeating in this process...(green,red) when green selected first and (red,green) when red selected first..since both the combinations are same..subtract 1 from the 6 ways..So, its 5

OpenStudy (anonymous):

i solve it as 1 1R AND 1Y=1C1*1C1=1 1R AND 1B=1C1*1C1=1 1G AND 1IY=1C1*1C1=1 1G AND 1B=1C1*1C1=1 1G AND 1R=1C1*1C1=1 SO TOTAL IS 5 WT DO U THINK

OpenStudy (anonymous):

Its correct and good with this problem as the numbers are small...but appears lengthy when the numbers are large..

OpenStudy (anonymous):

OK U R RIGHT

OpenStudy (anonymous):

Select... 1G in 1C1=1 way. The other box can be of any colour from the remaining three colours..So, you can select the second box in 3C1 = 3 ways... Similarly, do it for red too..But here, the second box should not be green as the combination is already selected above..So, the second box can be selected in 2C1=2ways.. Total ways=5

OpenStudy (anonymous):

THNX TO OPEN MY EYES

OpenStudy (anonymous):

i was little bit wrong ,,

OpenStudy (anonymous):

welcome : )